If possible, sole each sytem of equations. Use any method. If there is not a unique solution to a system, state a reason. 2x-8y=5 -6x+24y=15
Found 2 solutions by josgarithmetic, stanbon:
Answer by josgarithmetic(39617)

The slopes are the same.

Second equation when divide by 3 is -2x+8y=5. Multiply this by -1 and it becomes 2x-8y=-5.

The lines are different, but parallel. No solution.

Answer by stanbon(75887)

If possible, solve each sytem of equations. Use any method. If there is not a unique solution to a system, state a reason.
2x-8y=5
-6x+24y=15
-------
Multiply thru top equation by 3:
6x - 24y = 15
-6x+ 24y = 15
------
Add and solve for "y":
0 + 0 = 30
---
Contradiction:
---
Answer: No solution
========================
Cheers,
Stan H.